# On the Spectral Clustering of a Class of Multigrid Preconditioners

### v2

@@ Abstract
We consider an algebraic multigrid (AMG) scheme for the direct solution of complex-valued square linear systems based on a recursive $2\times 2$ block partitioning of the coefficient matrix and study the optimal choices of its components. In particular, we complement existing results that characterize optimal choices for a nonsymmetric-cycle methods by analyzing the spectral behavior of its symmetrized variant. We analyze the error propagation operator of the two-level symmetric-cycle method by calculating its invariant subspaces and its nonzero eigenvalues that govern the behavior of the error after a single cycle. We show that the error propagation operator can be studied separately for pairs of modes, working as bases of the invariant subspaces. The main result is an explicit choice of smoothing parameters that makes all the pairs of modes respond identically, forcing the nontrivial eigenvalues of the error propagation operator to collapse to a single, a priori known value, i.e., we give an explicit choice for the smoothing parameters so that the nonzero part of the spectrum of the error propagation operator collapses to a single, a priori known point on the real line. As a consequence, we give a closed-form formula for the inverse of a general square matrix that, to the best of our knowledge, is new in the literature. Additionally, the framework provides a clear, self-contained description of an ideal AMG W-cycle, offering a concrete target for the design of related schemes. We illustrate the theory with direct applications to general matrices and with analyses of representative matrices arising in numerical methods.
@@ 1 Introduction
are ubiquitous in scientific computing and modeling involving partial differential equations. Hence, the efficiency of solving eq. (1) is paramount for these applications.
Many of the most powerful solution techniques exploit some sort of hierarchy to approximate the discretized solution. Examples include multigrid methods [Hackbusch1985, Wesseling1992], domain decomposition methods with coarse correction [dolean2015introduction], hierarchical direct solvers [hackbusch2015hierarchical, Martinsson2015], and many parallel-in-time methods [gander2015FiftyYrs]. Multigrid methods in particular construct such hierarchies by combining smoothing with coarse representations of the problem on each level of the hierarchy so that at different levels the method focuses on different components of the error. Their success then depends on the interaction of the method’s components: the smoother, the smoothing parameters (i.e., damping and steps), the operators that transfer information between the levels of the hierarchy (i.e., restrictions and prolongations) and the coarse problems that represent the global behavior of the solution. Despite the wide use of multigrid methods, their processes are, in many cases, not understood in full detail, especially for non-Hermitian problems. In some applications, one must often rely on case-dependent analysis or numerical experimentation to obtain insight into the method’s mechanism for their problem [lucerolorca2024optimization, hemker2004fourier].
@@
This question is practically relevant – the vast majority of problems of the form of eq. (1) are solved only approximately through iterative methods. Many if not most of these iterative methods are formulated as an economical modification or approximation of direct solvers. This shows that a new direct solver opens the doors for new iterative methods, for example see [MacLachlanManteuffelMcCormick2006, LuceroLorca2025]. This remains true even in the case where the direct solver itself may suffer from numerical issues. Multigrid literature shows the condition on the regularity of the smoother is a natural one, frequently requiring the smoother to be a convergent method on its own[Wesseling1992]; we comment on the utility of the symmetry of the cycle at the end of this section.
@@
In the language of the multigrid community, we take the block-Jacobi smoother induced by the blocking of $L$, the ideal restriction and prolongation operators used in the nonsymmetric-cycle direct method and the inherited coarse problem and then show that with these choices we can construct a direct symmetric W-cycle method.
@@
We begin with a spectral analysis of the error propagation operator based on the smoothing parameters, under the assumption of diagonalizability of the Schur complement matrix pencil and show that the diagonalizability of this pencil ensures a decomposition of $\mathbb{C}^{N}$ into a direct sum of low-dimensional invariant subspaces of the error propagation operator. Analyzing its spectral properties can then be done independently in each of these subspaces, making the calculations tractable. We show that within each of these subspaces the error propagation operator has only a single nonzero eigenvalue, which depends on an eigenvalue of the pencil and on the smoothing parameters. Next, we give explicit formulas for the values of the smoothing parameters (depending on the number of pre- and post-smoothing steps) so that the nonzero eigenvalues become identical across all of the invariant subspaces, i.e., collapsing the spectrum of the error propagation operator into two a priori known points, one of which is zero. The derivation of this result is purely algebraic and hence independent of the origins of $L$, making the result relevant to a wide set of problems of the form of eq. (1). We then show several implications for making the two-level method direct and point out that it can be called recursively, obtaining a multilevel direct method.
@@
If a direct multigrid (or other) method is modified or approximated, we usually use Krylov subspace methods (KSM) to ensure or enhance its convergence, i.e., we use it as a preconditioner for KSM. Classical choices include the generalized minimal residual (GMRES) method or, in the Hermitian positive definite (HPD) setting, the conjugate gradient (CG) method (see [liesen2013krylov]). A single application of the modified/approximated multigrid method is carried out at each iteration of the KSM, thus building a more suitable Krylov subspace for the approximation of the solution compared to the unpreconditioned KSM. Importantly, if the multilevel method uses a nonsymmetric cycle and we use it as a preconditioner for an HPD problem, then CG can no longer be used, leading to loss of efficiency. This further motivates formulating a symmetric cycle, positive definite direct method (for cases when $L$ is HPD) so that even after its (suitable) approximation we can use it as a preconditioner for CG. We also note that in some cases, e.g., for CG, the convergence can be understood and predicted based on the eigenproperties of the preconditioned system. The spectral analysis of the error propagation operator of the two-level method directly yields analogous results for the preconditioned system when the two-level method is used as a preconditioner.
@@
The rest of this paper is structured as follows: we first refer to the existing results in the field in Section 1.1 and then recall some background material in Section 2. In Section 3 we develop the spectral properties of the error propagation operator of the two-level method and in Section 4 we use these result to derive the optimal choice of the smoothing parameters for the eigenvalue clustering of the error propagation operator and show how to obtain a direct method in Section 5. We provide numerical insights into the use of the method in Section 6.
@@ 1.1 Related work
Multigrid methods have a long history, beginning with analyses of relaxation for elliptic difference equations [Fedorenko1962, Fedorenko1964, Bakhvalov1966, Brandt1977, Hackbusch1985, Wesseling1992, BriggsHensonMcCormick2000], quickly accompanied with their algebraic counterparts – the algebraic multigrid – that evolved from the coarse and fine splitting framework, also called CF-splitting, introduced and developed in [Brandt1986, BrandtMcCormickRuge1985, RugeStueben1987]. This viewpoint leads to a direct two-level method approach that is generally impractical but serves as the theoretical foundation for many AMG algorithms, see Algorithm 1 in Section 2.
The nonsymmetric AMG literature offers several approaches for adapting this two-level method to challenging problems. A natural approach is to approximate ideal restriction and prolongation operators under fixed sparsity patterns with constrained optimization techniques [OlsonSchroderTuminaro2011, WiesnerTuminaroWallGee2014]. Petrov–Galerkin smoothed aggregation introduces transfer operators adapted to nonsymmetric systems [SalaTuminaro2008, BrezinaManteuffelMcCormickRugeSanders2010]. AIR and nAIR algorithms [ManteuffelRugeSouthworth2018] develop further the reduction-based AMG, showing that local approximations of ideal restriction combined with F-relaxation can yield a robust two-level method for general nonsymmetric problems. There are also adaptive strategies [MacLachlanManteuffelMcCormick2006, BrannickFrommerKahlMacLachlanZikatanov2010] or nonsymmetric methods based on approximate ideal restriction [ManteuffelMuenzenmaierRugeSouthworth2019].
Local Fourier Analysis (LFA) provides a tool for analyzing the multigrid convergence and smoothing behavior on structured grids, complementary to the analysis present in the above cited work. Classical LFA results [BuzbeeGolubNielson70, Brandt1977, BriggsHensonMcCormick2000] analyze the Fourier symbol of the operator and the relaxation scheme to understand the smoothing efficiency and the asymptotic two-grid contraction factor. Variants of this approach have been applied to discontinuous Galerkin methods and related discretizations, including settings where exact coarse-grid elimination leads to a direct two-level solver [LuceroLorcaGander2022].
For the HPD setting, the question of an optimal algorithm has been considered in [FalgoutVassilevski2004, XuZhang2018], including the energy-minimizing interpolation.
@@ 2 Background
with $L_{ff}\equiv A\in\mathbb{C}^{(N-n)\times(N-n)}$ and $L_{cc}\equiv D\in\mathbb{C}^{n\times n}$ being invertible. With this blocking, one can write the action of $L^{-1}$ as a two-level scheme, simply due to the Schur complement formulation of the LDU factorization of $L$, see [MacLachlanManteuffelMcCormick2006, LuceroLorca2025, FoersterSteubenTrottenberg1981, RiesTrottenbergWinter1983, ChanWan2000]. We recall this scheme in Algorithm 1, using the ideal prolongation, restriction and smoothing operators
@@
We consider a symmetric adaptation of Algorithm 1, with $m$ pre- and post-smoothing steps, each relaxed by a smoothing parameter $\alpha_{i}\in\mathbb{C}$, as summarized in Algorithm 2. The symmetric method is characterized by the prolongation, restriction and smoothing operators $P,R,S^{-1}$ and the smoothing parameters $\alpha_{1},\dotsc,\alpha_{m}$ – we automatically use the algebraically induced coarse operator $M_{0}:=RLP$.
@@
which then implies $M_{0}=D-CA^{-1}B$, i.e., the coarse space operator corresponds to the Schur complement of $A$ in $L$, same as in Algorithm 1.
We note that eq. (3) corresponds to taking $S^{-1}$ as the block Jacobi smoother, as opposed to using just its leading principal submatrix in Algorithm 1 (Jacobi-type smoothers are a common choice in the wider multigrid community); we adapted the prolongation and restriction operators directly from eq. (2) with $\star=-CA^{-1}$, ensuring that if $L$ is HPD then so are $M^{-1}$ and $M_{0}^{-1}$.
To have a concrete method, it remains to specify the smoothing parameters $\alpha_{1},\dotsc\alpha_{m}$, a problem treated in detail in Section 4. We conclude this section by recalling the error propagation operators of Algorithm 2. These are commonly used in multigrid analysis (see, e.g., [MacLachlanManteuffelMcCormick2006, eq. (2) and below]).
@@ 3 Spectral properties of $E$
Consider $E^{(s)}(\alpha)$ for some $\alpha\neq 0$, let $\mu:=(\lambda+\alpha-1)/\alpha$ and assume $\mu\neq 0$. The characteristic polynomial of $E^{(s)}(\alpha)$ is given by
@@
The spectrum of $E^{(c)}$ is equal to $\{0,1\}$ and so it seems intuitive that the spectrum of $E$ should also be parametrized by that of $T$. Next, we show that this intuition carries through – if $T$ is diagonalizable, then we can essentially block-diagonalize both $E^{(s)}(\alpha)$ and $E^{(c)}$ in the same basis [^2] We note that an analogous calculation and intuition also carries through if we work with the “complementary” matrix $D^{-1}CA^{-1}B$. Algebraically, all of the derivations that follow can be based on $D^{-1}CA^{-1}B$ with some minor adjustments.. First, we showcase a simplified version of our argument for the case n=Nn=N/2n=N-n=N/2 where the blocks $A,\dotsc,D$ are square and assuming $T$ is regular. On one hand, this leads to convenient simplifications that lay out the structure of the operators clearly; on the other hand, these assumptions are mostly unrealistic as we almost always have n>Nnn>N-n. We treat the general case in analogous (though more tedious) fashion afterwards. We note that the case n<Nnn<N-n can be handled “by symmetry” – in the calculation above, we could use the other Schur complement formulation for the determinant, featuring the matrix $D^{-1}CA^{-1}B$ instead of $T$. This would lead to calculations that are analogous to the ones we carry out below, replacing the eigenpairs of $T$ with those of $D^{-1}CA^{-1}B$ and exchanging the blocks $D^{-1}C$ and $A^{-1}B$ wherever necessary.
@@ 3.1 Invariant subspaces – the simplified case
In this section we compute the invariant subspaces of the error propagation operator assuming that $n=N/2$, where $n$ is the size of the coarse space and $N$ the size of the fine space, and $T\in\mathbb{C}^{N/2\times N/2}$ is regular and diagonalizable. Hence, we have a full set of $N/2$ eigenpairs $(\lambda_{1},\mathbf{w}_{1}),\dotsc,(\lambda_{N/2},\mathbf{w}_{N/2})$, where all eigenvalues are nonzero. Let the eigendecomposition of $T$ be $W\Lambda W^{-1}$. Define the matrix $\tilde{V}$ as
@@
Since $T$ is regular so is $C$ and $\tilde{V}$. We can transform $E^{(s)}(\alpha)$ and $E^{(c)}$ into the basis given by the columns of $\tilde{V}$. A direct calculation gives
@@
$V^{-1}E^{(c)}V=\frac{1}{2}\begin{bmatrix}\begin{bmatrix}1+\sqrt{\lambda_{1}}&1-\sqrt{\lambda_{1}}\\ 1+\sqrt{\lambda_{1}}&1-\sqrt{\lambda_{1}}\end{bmatrix}\\ &\ddots\\ &&\begin{bmatrix}1+\sqrt{\lambda_{N/2}}&1-\sqrt{\lambda_{N/2}}\\ 1+\sqrt{\lambda_{N/2}}&1-\sqrt{\lambda_{N/2}}\end{bmatrix}\end{bmatrix}$.
In other words, we have found $N/2$ two-dimensional invariant subspaces $\mathcal{V}_{\lambda_{1}},\dotsc,\mathcal{V}_{\lambda_{N/2}}$ of the error propagation operators $E^{(c)},E^{(s)}(\alpha)$ and thus also of any of their compositions. Therefore, we can analyze the global error propagation operator $E(\alpha_{1},\dotsc,\alpha_{m})$ defined in eq. (5) on the subspace $\mathcal{V}:=\mathcal{V}_{\lambda_{1}}\oplus\dotsc\oplus\mathcal{V}_{\lambda_{N/2}}$ (with the basis $V$) using $2\times 2$ matrices, allowing for a direct calculation. In the next section we derive analogous results in the general case but first, we briefly comment on the symbol $\sqrt{\Lambda}\equiv\mathrm{diag}(\sqrt{\lambda_{1}},\dotsc,\sqrt{\lambda_{N/2}})$ in the following remark.

### v3

@@ Abstract
We consider an algebraic multigrid (AMG) scheme for the direct solution of complex-valued square linear systems based on a recursive $2\times 2$ block partitioning of the coefficient matrix and study the optimal choices of its components. In particular, we complement existing results that characterize the optimal choices for a nonsymmetric cycle method by analyzing the spectral behavior of its symmetric cycle variant. We analyze the error propagation operator of a specific two-level symmetric cycle method by calculating its invariant subspaces and its nonzero eigenvalues that influence the behavior of the error after a single cycle. We show that the error propagation operator can be studied separately for pairs of modes, working as bases of the invariant subspaces. The main result is an explicit choice of smoothing parameters that makes all the pairs of modes respond identically, forcing the nontrivial eigenvalues of the error propagation operator to collapse to a single, a priori known value. As a consequence, we give a new closed form formula for the inverse of a general square matrix. Additionally, the framework provides a clear, self-contained description of an ideal AMG K-cycle, offering a concrete target for the design of related schemes. We illustrate the theory with direct applications to general matrices and with analyses of representative matrices arising in numerical methods.
@@ 1 Introduction
are ubiquitous in scientific computing and modeling involving partial differential equations. Hence, the efficiency of solving eq. (1) is critical for these applications.
Many of the most powerful solution techniques exploit some sort of hierarchy to approximate the discretized solution. Examples include multigrid methods [Hackbusch1985, Wesseling1992], domain decomposition methods with coarse correction [dolean2015introduction], hierarchical direct solvers [hackbusch2015hierarchical, Martinsson2015], and many parallel-in-time methods [gander2015FiftyYrs]. Multigrid methods in particular construct such hierarchies by combining smoothing with coarse representations of the problem on each level of the hierarchy so that at different levels the method focuses on different components of the error. Their success then depends on the interaction of the method’s components: the smoother, the smoothing parameters (i.e., damping and number of steps), the operators that transfer information between the levels of the hierarchy (i.e., restrictions and prolongations) and the coarse problems that represent the global behavior of the solution. Despite the wide use of multigrid methods, in many cases their processes are not understood in full, especially for non-Hermitian problems. In some applications, one must often rely on case-dependent analyses or numerical experimentation to obtain insight into the method’s mechanism for a specific problem [lucerolorca2024optimization, hemker2004fourier].
@@
This question is practically relevant – the vast majority of problems of the form of eq. (1) are solved only approximately through iterative methods. Many of these iterative methods are formulated as an economical modification or approximation of direct solvers. This shows that a new direct solver opens the doors for new iterative methods, for example see [MacLachlanManteuffelMcCormick2006, LuceroLorca2025]. This remains true even in the case where the direct solver itself may suffer from numerical issues. The multigrid literature shows that the condition on the regularity of the smoother is a natural one, frequently requiring the smoother to be a convergent method on its own [Wesseling1992]; we comment on the utility of the symmetry of the cycle at the end of this section.
@@
In the language of the multigrid community, we take the block-Jacobi smoother induced by the blocking of $L$, the ideal restriction and prolongation operators used in the nonsymmetric cycle direct method and the inherited coarse problem and then show that with these choices we can construct a direct method that can be expressed both as a direct formula and as a symmetric K-cycle method with the coarse correction called twice recursively, i.e. a Krylov W-cycle.
@@
We begin with a spectral analysis of the error propagation operator based on the smoothing parameters, under the assumption of diagonalizability of the Schur complement matrix pencil and show that the diagonalizability of this pencil ensures a decomposition of $\mathbb{C}^{N}$ into a direct sum of low-dimensional invariant subspaces of the error propagation operator. Analyzing its spectral properties can then be done independently in each of these subspaces, making the calculations tractable. We show that within each of these subspaces the error propagation operator has only a single nonzero eigenvalue, which depends on the spectrum of the pencil and on the smoothing parameters. Next, we give explicit formulas for the values of the smoothing parameters (depending on the number of pre- and post-smoothing steps) so that the nonzero eigenvalues become identical across all of the invariant subspaces, i.e., collapsing the spectrum of the error propagation operator into two a priori known points, one of which is zero. The derivation of this result is purely algebraic and hence independent of the origins of $L$, making the result relevant to a wide set of problems of the form of eq. (1). We then show several implications for making the two-level method direct and point out that it can be called recursively, obtaining a multilevel direct method.
@@
If a direct multigrid (or other) method is modified or approximated, we usually use Krylov subspace methods (KSM) to ensure or enhance its convergence, i.e., we use the approximated direct method as a preconditioner for KSM. Classical choices include the generalized minimal residual (GMRES) method or, in the Hermitian positive definite (HPD) setting, the conjugate gradient (CG) method. A single application of the modified/approximated multigrid method is carried out at each iteration of the KSM, thus building a more suitable Krylov subspace for the approximation of the solution compared to the unpreconditioned KSM. Importantly, if the multilevel method uses a nonsymmetric cycle and we use it as a preconditioner for an HPD problem, then CG can no longer be used, leading to loss of efficiency. This further motivates formulating a symmetric cycle, positive definite direct method (for cases when $L$ is HPD) so that even after its (suitable) approximation we can use it as a preconditioner for CG. We also note that in some cases, e.g., for CG, the convergence can be understood and predicted based solely on the eigenproperties of the preconditioned system. The spectral analysis of the error propagation operator of the two-level method directly yields analogous results for the preconditioned system when the two-level method is used as a preconditioner. We refer the reader to [liesen2013krylov] for further references on KSM and their convergence.
@@
The rest of this paper is structured as follows: in the remainder of this chapter we refer to the existing results in the field. We recall background material in Section 2. In Section 3 we develop the spectral properties of the error propagation operator of the two-level method. In Section 4 we use these result to derive the optimal choice of the smoothing parameters for the eigenvalue clustering of the error propagation operator. We show how to obtain a direct method in Section 5. We provide numerical insights into the use of the method in Section 6.
@@ 1.1 Related work
Multigrid methods have a long history, beginning with analyses of relaxation for elliptic difference equations [Fedorenko1962, Fedorenko1964, Bakhvalov1966, Brandt1977, Hackbusch1985, Wesseling1992, BriggsHensonMcCormick2000], then quickly followed by their algebraic counterparts – the algebraic multigrid methods (AMG) – introduced and developed in [Brandt1986, BrandtMcCormickRuge1985, RugeStueben1987]. These AMG evolved from the coarse and fine splitting framework, also called CF-splitting. This viewpoint leads to a direct two-level method approach that serves as the theoretical foundation for many AMG algorithms, see Algorithm 1 in Section 2.
The nonsymmetric AMG literature offers several approaches for adapting this two-level method to challenging problems. A natural approach is to approximate the ideal restriction and prolongation operators under fixed sparsity patterns with constrained optimization techniques [OlsonSchroderTuminaro2011, WiesnerTuminaroWallGee2014]. Petrov–Galerkin smoothed aggregation introduces transfer operators adapted to nonsymmetric systems [SalaTuminaro2008, BrezinaManteuffelMcCormickRugeSanders2010]. AIR and nAIR algorithms [ManteuffelRugeSouthworth2018] develop further the reduction-based AMG, showing that local approximations of ideal restriction combined with F-relaxation can yield a robust two-level method for general nonsymmetric problems. There are also adaptive strategies [MacLachlanManteuffelMcCormick2006, BrannickFrommerKahlMacLachlanZikatanov2010] or nonsymmetric methods based on an approximate ideal restriction [ManteuffelMuenzenmaierRugeSouthworth2019].
Local Fourier Analysis (LFA) provides a tool for analyzing the multigrid convergence and smoothing behavior on structured grids, complementary to the analysis in the above cited work. Classical LFA results [BuzbeeGolubNielson70, Brandt1977, BriggsHensonMcCormick2000] analyze the Fourier symbol of the operator and the relaxation scheme to understand the smoothing efficiency and the asymptotic two-grid contraction factor. Variants of this approach have been applied to discontinuous Galerkin methods and related discretizations, including settings where exact coarse-grid elimination leads to a direct two-level solver [LuceroLorcaGander2022].
In the HPD setting, the existence of an optimal algorithm has been considered in [FalgoutVassilevski2004, XuZhang2018], by means of an energy-minimizing interpolation.
@@ 2 Background
with $L_{ff}\equiv A\in\mathbb{C}^{(N-n)\times(N-n)}$ and $L_{cc}\equiv D\in\mathbb{C}^{n\times n}$ being invertible. With this blocking, one can write the action of $L^{-1}$ as a two-level scheme, due to the Schur complement formulation of the LDU factorization of $L$, see [MacLachlanManteuffelMcCormick2006, LuceroLorca2025, FoersterSteubenTrottenberg1981, RiesTrottenbergWinter1983, ChanWan2000]. We recall this scheme in Algorithm 1, using the ideal prolongation, restriction and smoothing operators
@@
We consider a symmetric cycle adaptation of Algorithm 1, with $m$ pre- and post-smoothing steps, each relaxed by a smoothing parameter $\alpha_{i}\in\mathbb{C}$, as summarized in Algorithm 2. The symmetric cycle method is characterized by the prolongation, restriction and smoothing operators $P,R,S^{-1}$ and the smoothing parameters $\alpha_{1},\dotsc,\alpha_{m}$ – we automatically use the algebraically induced coarse operator $M_{0}:=RLP$.
@@
which then implies $M_{0}=D-CA^{-1}B$, i.e., the coarse space operator corresponds to the Schur complement of $A$ in $L$, as it did in Algorithm 1.
We note that eq. (3) corresponds to taking $S^{-1}$ as the block Jacobi smoother, as opposed to using just its leading principal submatrix in Algorithm 1 (Jacobi-type smoothers are a common choice in the wider multigrid community); we adapted the prolongation and restriction operators directly from eq. (2) with $\star=-CA^{-1}$, ensuring that if $L$ is HPD, then so are $M^{-1}$ and $M_{0}^{-1}$.
It remains to specify the smoothing parameters $\alpha_{1},\dotsc\alpha_{m}$, a problem treated in detail in Section 4. We conclude this section by recalling the error propagation operators of Algorithm 2. These are commonly used in multigrid analysis (see, e.g., [MacLachlanManteuffelMcCormick2006, eq. (2) and below]).
@@ 3 Spectral analysis of $E$
Consider the characteristic polynomial of $E^{(s)}(\alpha)$ for some $\alpha\neq 0$ given by $\det(\lambda I-E^{(s)})$ and let $\mu:=(\lambda+\alpha-1)/\alpha$. Assuming $\mu\neq 0$ we can write
@@
The spectrum of $E^{(c)}$ is equal to $\{0,1\}$ and so one could hope that the spectrum of $E$ is also parametrized by that of $T$. Next, we show that this intuition carries through – if $T$ is diagonalizable, then we can essentially block-diagonalize both $E^{(s)}(\alpha)$ and $E^{(c)}$ in the same basis. First, we showcase a simplified version of our argument for the case n=Nn=N/2n=N-n=N/2 where the blocks $A,\dotsc,D$ are square and we assume $T$ is regular. On the one hand, this leads to convenient simplifications that lay out the structure of the operators clearly; on the other hand, these assumptions are mostly unrealistic as we almost always have n>Nnn>N-n. Afterwards, we treat the general case in analogous fashion. We note that the case n<Nnn<N-n can be handled by symmetry – in the calculation above, we could use the other Schur complement formulation for the determinant, featuring the matrix $D^{-1}CA^{-1}B$ instead of $T$. This would lead to calculations that are analogous to the ones we carry out below, replacing the eigenpairs of $T$ with those of $D^{-1}CA^{-1}B$ and exchanging the blocks $D^{-1}C$ and $A^{-1}B$ wherever necessary.
@@ 3.1 Invariant subspaces – the simplified case
In this section we compute the invariant subspaces of the error propagation operator assuming that $n=N/2$, where $n$ is the size of the coarse space and $N$ is the size of the original problem, and $T\in\mathbb{C}^{N/2\times N/2}$ is regular and diagonalizable. Hence, we have a full set of $N/2$ eigenpairs $(\lambda_{1},\mathbf{w}_{1}),\dotsc,(\lambda_{N/2},\mathbf{w}_{N/2})$, where all eigenvalues are nonzero. Let the eigendecomposition of $T$ be $W\Lambda W^{-1}$. Define the matrix $\tilde{V}$ as
@@
Since $T$ is regular so are $C$ and $\tilde{V}$. We can transform $E^{(s)}(\alpha)$ and $E^{(c)}$ into the basis given by the columns of $\tilde{V}$. A direct calculation gives
@@
[…]
In other words, we have found $N/2$ two-dimensional invariant subspaces $\mathcal{V}_{\lambda_{1}},\dotsc,\mathcal{V}_{\lambda_{N/2}}$ of the error propagation operators $E^{(c)},E^{(s)}(\alpha)$ and thus also of any of their compositions. Therefore, we can analyze the global error propagation operator $E(\alpha_{1},\dotsc,\alpha_{m})$ defined in eq. (5) on the subspace $\mathcal{V}:=\mathcal{V}_{\lambda_{1}}\oplus\dotsc\oplus\mathcal{V}_{\lambda_{N/2}}$ (with the basis $V$) using $2\times 2$ matrices, allowing for a direct calculation.
In the next section we derive analogous results in the general case but first, we briefly comment on the symbol $\sqrt{\Lambda}\equiv\mathrm{diag}(\sqrt{\lambda_{1}},\dotsc,\sqrt{\lambda_{N/2}})$ in the following remark.